# Plain Transformers Can be Powerful Graph Learners

### v2

@@ 5.3 Ablation Study on Proposed Designs
[…]
Further Study
We present further experimental studies and results in Appx. B.
@@ 6 Conclusion
Plain transformers are ill-suited for handling the unique challenges posed by graphs, such as the lack of canonical coordinates and permutation invariance. To obtain superior capacity and empirical performance, previous graph transformers (GTs) have introduced non-standard, sophisticated, and domain-specific architectural modifications. In this work, we demonstrate that plain transformers can be powerful graph learners via the proposed minimal, easy-to-adapt modifications. Our Powerful Plain Graph Transformers (PPGT) not only achieve competitive expressivity, but also demonstrate outstanding empirical performance on real-world graph benchmarks while maintaining the simplicity of plain Transformer architectures. Further exploration of empirical expressivity also unveils a potential direction for improving plain GTs. We consider this work an important first step toward reducing dissimilarities between GTs and transformers in other domains. This direction of investigation is crucial for building general multi-modal foundation models that integrate learning capabilities on graphs and potentially other irregular non-Euclidean geometric spaces.
@@
This work demonstrates that plain Transformers can be powerful graph learners. However, like most plain Transformers, PPGTs still require the $O(N^{2})$ computational complexity and might encounter issues on scaling to large-size inputs. A more detailed discussion of limitations and possible avenues to address them is provided in Appx. F.
@@ Appendix A Experimental Details
A.1 Bechmarking GNNs and Long Range Graph Benchmarks
@@
Table 5 provides a summary of the statistics and characteristics of datasets used in this paper. The first five datasets are from Dwivedi et al., 2022a , and the last two are from Dwivedi et al., 2022c . Readers are referred to Rampášek et al., (2022) for more details of the datasets.
Table 5: Overview of the graph learning datasets involved in this work Dwivedi et al., 2022a ; Dwivedi et al., 2022c ; Irwin et al., (2012); Hu et al., (2021).
@@
Dataset splits, random seed and parameter budgets
We conduct the experiments on the standard train/validation/test splits of the evaluated benchmarks, following previous works Rampášek et al., (2022); Ma et al., (2023). For each dataset, we execute 4 trials with different random seeds (0, 1, 2, 3) and report the mean performance and standard deviation. We follow the most commonly used parameter budgets: around 500k parameters for ZINC, PATTERN, CLUSTER, Peptides-func, and Peptides-struct; and around 100k parameters for SP-MNIST and SP-CIFAR10.
@@
Due to the limited time and computational resources, we did not perform an exhaustive search on the hyperparameters. We start with the hyperparameter setting of GRIT Ma et al., (2023) and perform minimal search to satisfy the commonly used parameter budgets.
The hyperparameters are presented in Table 6 and Table 7. In the tables, S.D. stands for stochastic depth Huang et al., (2016), a.k.a., drop-path, in which we treat a graph as one example in stochastic depth.
@@
Table 6: Hyperparameters for five datasets from Benchmarking GNNs Dwivedi et al., 2022a
@@
Table 7: Hyperparameters for two datasets from the Long-range Graph Benchmark Dwivedi et al., 2022c
Hyperparameter Peptides-func Peptides-struct
@@
MLP-dim 128 128
# FFN 1 1
@@
S.D. 0.1 0.1
Attn. dropout 0.2 0.1
@@
Graph Pooling sum mean
Norm BN -
# layers 3 2
PE-dim 32 24
@@
Batch size 32 32
Learn. rate 7e-4 7e-4
# epochs 400 250
# warmup 10 10
Weight decay 1e-5 1e-5
# Parameters 509K 488K
A.2 BREC: Empirical Expressivity Benchmark
BREC Wang and Zhang, (2024) is an empirical expressivity datasets, consisting of four major categories of graphs: Basic, Regular, Extension, and CFI. Basic graphs include $60$ pairs of simple 1-WL-indistinguishable graphs. Regular graphs include $140$ pairs of regular graphs from four types of subcategories: $50$ pairs of simple regular graphs, $50$ pairs of strongly regular graphs, $20$ pairs of 4-vertex condition graphs and $20$ pairs of distance regular graphs. Extension graphs include $100$ pairs of special graphs that arise when comparing four kinds of GNN extensions Papp and Wattenhofer, (2022). CFI graphs include $100$ pairs of graphs generated by CFI methods Cai et al., (1992). All pairs of graphs are 1-WL-indistinguishable.
We follow the standard training pipelines from BREC: pairwise contrastive training and evaluate process. No specific parameter budget required on BREC dataset.
Hyperparameter
The hyperparameter setting for PPGT on BREC can be found in Table 8.
For the experimental setup of I2-GNN + PPGT, we construct a subgraph for each edge within the 4-hop neighborhood and use RRWP as the node-attributes. We employ a 6-layer GIN with BN on each subgraph independently to get its representation. The subgraph representations are fed to PPGT as the additional edge-attributes. The I2-GNN and PPGT are trained end-to-end together.
Table 8: Hyperparameters for BREC Wang and Zhang, (2024)
@@ A.2 BREC: Empirical Expressivity Benchmark
A.3 PCQM4Mv2 from OGB Large-Scale Challenge
PCQM4Mv2 is a large-scale quantum chemistry graph dataset benchmark, containing over 3.7M graphs, proposed from OGB Large-Scale Challenge (OGB-LSC) Hu et al., (2021). The statistics of the dataset can be found in Table 5.
A.4 Optimizer and Learning Rate Scheduler
Following most plain Transformers in other domains, we use AdamW Loshchilov and Hutter, (2019) as the optimizer and the Cosine Annealing Learning Rate scheduler Loshchilov and Hutter, (2017) with linear warm up.
A.5 Baselines Information
For benchmarks from BenchmarkingGNN Dwivedi et al., 2022a .
  • Tyical Message-passing Networks(MPNNs): GCN Kipf and Welling, (2017), GIN Xu et al., (2019), GAT Veličković et al., (2018), GatedGCN Bresson and Laurent, (2018), PNA Corso et al., (2020);
  • GNNs going beyond MPNNs: CRaW1 Tönshoff et al., (2023), GIN-AK+ Zhao et al., (2022), DGN Beani et al., (2021), CKGCN Ma et al., (2024),
  • Graph Transformers SAN Kreuzer et al., (2021), K-Subgraph SAT Chen et al., (2022), EGT Hussain et al., (2022), Graphormer-GD Zhang et al., 2023b , GPS Rampášek et al., (2022), GMLP-Mixer He et al., (2023), GRIT Ma et al., (2023).
For BREC Wang and Zhang, (2024):
  • Subgraph GNNs: SUN Frasca et al., (2022), SSWL+ Zhang et al., 2023a , I2-GNN Huang et al., (2023)
  • K-WL/K-FWL GNNs: PPGN Maron et al., (2019), 2-DRFWL(2) Zhou et al., (2023), 3-DRFWL(2) Zhou et al., (2023)
  • K-FWL+SubgraphGNNs: N2GNN Feng et al., (2023)
  • GD-WL GNNs Graphormer Ying et al., (2021), EPNN Zhang et al., 2024b , CKGConv Ma et al., (2024), GRIT Ma et al., (2023).
For OGB-LSC Hu et al., (2021):
MPNNs (GCN Kipf and Welling, (2017), GIN Xu et al., (2019) with/without virtual nodes) as well as several Graph Transformers (GRPE Park et al., (2022), Graphormer Ying et al., (2021), TokenGT Kim et al., (2022) and GraphGPS Rampášek et al., (2022)).
@@ B.1 Sensitivity Study: Impact of the Number of Bases $S$ in SPE on Empirical Expressivity
Table 9: Sensitivity study on $S$ in Sinusoidal encoding for RRWP on CFI graphs from BREC Wang and Zhang, (2024).
@@ B.2 Case Study Comparing BN, RMSN and AdaRMSN
Figure 5: (Case Study of AdaRMSN) Visualization of Input and Pred data points [(1). Input; (2) Predictions w/ BN; (3) Predictions w/ RMSN; (4) Predictions w/ AdaRMSN]. RMSN is ineffective in preserving magnitude information, whereas both BN and AdaRMSN successfully maintain the crucial magnitude information of the data point
@@
To better understand the advantages of AdaRMSN over BatchNorm (BN), we conducted a sensitivity study comparing normalization techniques w.r.t. batch sizes, based on PPGT on the ZINC dataset The study (as shown in Fig. 6) showcases the better stability of AdaRMSN compared to the choice of BN used in many previous GTs Kreuzer et al., (2021); Rampášek et al., (2022); Ma et al., (2023).
@@
B.4 Runtime and GPU Comsumption
We provide a basic statistics on runtime and gpu comsumption of PPGT, comparing to GRIT (as shown in Table 10).
@@ C.1 Attention in Graph Transformers
In SAN Kreuzer et al., (2021) and GRIT Ma et al., (2023), the attention mechanism is implemented with the sparse operations in PyTorch Geometric due to the complicated attention mechanism. For PPGTs, we provide two versions, one based on the sparse operations and another one based on the dense operation with padding. The latter one is typically faster for larger-scale graphs but consumes more memory.
@@ C.2 Initializations of Parameters
Following most recent plain Transformers, we initialize the weights of linear layers in the backbone as well as the prediction heads with truncated normalization with standard deviation $\sigma=0.02$.
For the stems, we utilize Kaiming uniform initialization He et al., (2015) with $a=0$ for hidden layers in MLPs and $a=1$ for output layers in MLPs or standalone linear layers.
@@ C.3 Injection of Degree and Graph Order
For superpixel datasets (CIFAR10, MNIST), we also include the location of pixel into graph PE.
@@ D.1 GRIT’s Attention
The attention mechanism in GRIT Ma et al., (2023) adopts the conditional MLP Perez et al., (2018), involves linear projections, elementwise multiplications, and an uncommon non-linearity in the form of a signed-square-root:
@@
where $\sigma$ is a non-linear activation (ReLU by default); $\mathbf{W}_{\text{Q}},\mathbf{W}_{\text{K}},\mathbf{W}_{\text{Ew}},\mathbf{W}_{\text{Eb}}\in\mathbb{R}^{d^{\prime}\times d}$, $\mathbf{W}_{\text{A}}\in\mathbb{R}^{1\times d^{\prime}}$, $\mathbf{W}_{\text{V}}\in\mathbb{R}^{d\times d}$ and $\mathbf{W}_{\text{Ev}}\in\mathbb{R}^{d\times d^{\prime}}$ are learnable weight matrices; $\odot$ indicates elementwise multiplication; and $\rho(\mathbf{x}):=(\text{ReLU}(\mathbf{x}))^{1/2}-(\text{ReLU}(-\mathbf{x}))^{1/2}$ is the signed-square-root. Note that $\mathbf{e}_{i,j}$ here corresponds to $\mathbf{p}_{ij}$ in our PPGT but it requires update within the attention mechanism. According to Ma et al., (2023), the signed-square-root $\rho$ is necessary to stabilize the training.

### v3

@@ 5.4 Ablation Study on Proposed Designs
5.5 Further Study
In Appx. B, we present additional experimental analyses, including: (B.1) a sensitivity study on the number of bases $S$ in SPE, demonstrating the impact of SPE on empirical expressivity; (B.2) a case study of different normalization layers, showing that RMSN discards magnitude information whereas AdaRMSN preserves it; (B.3) a batch-size sensitivity analysis of AdaRMSN, demonstrating its robustness to batch size, in contrast to BN; and (B.4) a comparison of runtime and GPU memory consumption between PPGT and GRIT.
@@ 6 Conclusion
Plain Transformers are ill-suited for handling the unique challenges posed by graphs, such as the lack of canonical coordinates and permutation invariance. To obtain superior capacity and empirical performance, previous graph Transformers (GTs) have introduced non-standard, sophisticated, and domain-specific architectural modifications. In this work, we demonstrate that plain Transformers can be powerful graph learners via the proposed minimal, easy-to-adapt modifications. Our Powerful Plain Graph Transformers (PPGTs) not only achieve competitive expressivity, but also demonstrate strong empirical performance on real-world graph benchmarks while maintaining the simplicity of plain Transformer architectures. Further exploration of empirical expressivity also unveils a potential direction for improving plain GTs.
We consider this work an important first step toward reducing dissimilarities between GTs and Transformers in other domains. This potentially eases the design of general multi-modal foundation models that integrate learning capabilities on graphs and potentially other irregular non-Euclidean geometric spaces.
Beyond graph domains, the insights from our work may also benefit Transformer architectures in other domains, broadening their applicability and impact.
@@
This work demonstrates that plain Transformers can serve as powerful graph learners without external graph models (e.g., MPNNs). However, like most vanilla Transformer architectures, PPGTs incur an $\mathcal{O}(N^{2})$ computational complexity, which limits their direct applicability to large-size input graphs. To mitigate this issue, graph sampling techniques can be employed to handle large inputs more efficiently. A more detailed discussion of these limitations, along with potential directions for addressing them, is provided in Appx. F.
@@ Appendix A Experimental Details
A.1 Benchmarking GNNs and Long Range Graph Benchmarks
@@
Table 6 provides a summary of the statistics and characteristics of the datasets used in this paper. The first five datasets are from Dwivedi et al. (2022a), and the last two are from Dwivedi et al. (2022c). Readers are referred to Rampášek et al. (2022) for more details of the datasets.
Table 6: Overview of the graph learning datasets involved in this work Dwivedi et al. (2022a; c); Irwin et al. (2012); Hu et al. (2021).
@@
PascalVoc-SP 11,355 479.4 2,710.5 No Inductive Node 21-class classif. Marco F1
@@
OGBN-ArXiv 1 169,343 1,116,243 Yes Transductive Node 40-class classif. Accuracy
Dataset splits, random seed, and parameter budgets
We conduct the experiments on the standard train/validation/test splits of the evaluated benchmarks, following previous works Rampášek et al. (2022); Ma et al. (2023). For each dataset, we execute 4 trials with different random seeds (0, 1, 2, 3) and report the mean performance and standard deviation. We follow the most commonly used parameter budgets: around 500k parameters for ZINC, PATTERN, CLUSTER, Peptides-func, and Peptides-struct; and around 100k parameters for SP-MNIST and SP-CIFAR10.
@@
Due to the limited time and computational resources, we did not perform an exhaustive search on the hyperparameters. We start with the hyperparameter setting of GRIT Ma et al. (2023) and perform minimal search to satisfy the commonly used parameter budgets.
The hyperparameters are presented in Table 7 and Table 8. In the tables, S.D. stands for stochastic depth Huang et al. (2016), a.k.a., drop-path, in which we treat a graph as one example in stochastic depth.
@@
Table 7: Hyperparameters for five datasets from Benchmarking GNNs Dwivedi et al. (2022a)
@@
Table 8: Hyperparameters for two datasets from the Long-range Graph Benchmark Dwivedi et al. (2022c)
Hyperparameter Peptides-func Peptides-struct PascalVoc-SP
@@
MLP-dim 128 128 128
Edge-dim 64 64 64
# FFN 1 1 1
FFN expansion 2 2 2
# bases 3 3 3
PE-dim 32 24 16
Backbone
# blocks 5 5 12
Dim 96 96 64
FFN expansion 2 2 2
# attn. heads 16 8 8
S.D. 0.1 0.1 0.2
Attn. dropout 0.2 0.1 0.3
Pred. Head
Graph Pooling sum mean -
Norm BN - -
# layers 3 2 2
Training
Batch size 32 32 16
Learn. rate 7e-4 7e-4 1e-3
# epochs 400 250 200
# warmup 10 10 10
Weight decay 1e-5 1e-5 1e-2
# Parameters 509K 488K 492K
A.2 BREC: Empirical Expressivity Benchmark
BREC Wang and Zhang (2024) is an empirical expressivity dataset, consisting of four major categories of graphs: Basic, Regular, Extension, and CFI. Basic graphs include $60$ pairs of simple 1-WL-indistinguishable graphs. Regular graphs include $140$ pairs of regular graphs from four types of subcategories: $50$ pairs of simple regular graphs, $50$ pairs of strongly regular graphs, $20$ pairs of 4-vertex condition graphs and $20$ pairs of distance regular graphs. Extension graphs include $100$ pairs of special graphs that arise when comparing four kinds of GNN extensions Papp and Wattenhofer (2022). CFI graphs include $100$ pairs of graphs generated by CFI methods Cai et al. (1992). All pairs of graphs are 1-WL-indistinguishable.
We follow the standard training pipelines from BREC: pairwise contrastive training and evaluating process. No specific parameter budget required on the BREC dataset.
Hyperparameter
The hyperparameter setting for PPGT on BREC can be found in Table 10.
For the experimental setup of I2-GNN + PPGT, we construct a subgraph for each edge within the 4-hop neighborhood and use RRWP as the node-attributes. We employ a 6-layer GIN with BN on each subgraph independently to get its representation. The subgraph representations are fed to PPGT as additional edge-attributes. The I2-GNN and PPGT are trained end-to-end together.
Table 9: Hyperparameters for PCQM4Mv2 Hu et al. (2021) and OGBN-ArXiv Hu et al. (2020).
Hyperparameter PCQM4Mv2 OGBN-ArXiv
PE Stem
MLP-dim 512 192
Edge-dim 128 96
# FFN 8 2
@@ A.2 BREC: Empirical Expressivity Benchmark
S.D. 0.2 0.5
Attn. dropout 0.2 0.5
Node2Subgraph
Node Masking - 0.5
Max Size of Graph - 100
@@
Graph Pooling sum class attention
Norm - -
# layers 3 1
@@
Batch size 2048 128
Learn. rate 1e-3 2e-3
# epochs 500 300
# warmup 50 10
Weight decay 1e-3 5e-2
# parameters 17.6M 3.74M
A.3 PCQM4Mv2 from OGB Large-Scale Challenge
PCQM4Mv2 is a large-scale quantum chemistry graph dataset benchmark, containing over 3.7M graphs, proposed from OGB Large-Scale Challenge (OGB-LSC) Hu et al. (2021). The statistics of the dataset can be found in Table 6.
A.4 OGBN-ArXiv from OGB Benchmark
OGBN-ArXiv is a large-scale graph benchmark featuring a transductive node classification task from Open Graph Benchmark Hu et al. (2020). Unlike the other graph datasets used in this work, OGBN-ArXiv comprises a single large-scale graph containing over 169,000 nodes. This benchmark emphasizes the graph models’ scalability to large input graphs. The statistics of the dataset can be found in Table 6.
For PPGT, we convert the transductive node-classification task on OGBN-ArXiv into a graph-classification setting via the Node2Subgraph transformation. In this framework, we employ the class-attention mechanism Touvron et al. (2021b) as the output head. As OGBN-ArXiv is a directed graph, we compute dual-direction RRWP features, i.e., random walks along both in-edge and out-edge directions.
Table 10: Hyperparameters for BREC Wang and Zhang (2024)
@@ A.4 OGBN-ArXiv from OGB Benchmark
A.5 Optimizer and Learning Rate Scheduler
Following most plain Transformers in other domains, we use AdamW Loshchilov and Hutter (2019) as the optimizer and the Cosine Annealing Learning Rate scheduler Loshchilov and Hutter (2017) with linear warm up.
A.6 Baselines Information
For benchmarks from BenchmarkingGNN Dwivedi et al. (2022a).
  • Tyical Message-passing Networks(MPNNs): GCN Kipf and Welling (2017), GIN Xu et al. (2019), GAT Veličković et al. (2018), GatedGCN Bresson and Laurent (2018), PNA Corso et al. (2020);
  • GNNs going beyond MPNNs: CRaW1 Tönshoff et al. (2023), GIN-AK+ Zhao et al. (2022), DGN Beani et al. (2021), CKGCN Ma et al. (2024),
  • Graph Transformers SAN Kreuzer et al. (2021), K-Subgraph SAT Chen et al. (2022), EGT Hussain et al. (2022), Graphormer-GD Zhang et al. (2023b), GPS Rampášek et al. (2022), GMLP-Mixer He et al. (2023), GRIT Ma et al. (2023).
For BREC Wang and Zhang (2024):
  • Subgraph GNNs: SUN Frasca et al. (2022), SSWL+ Zhang et al. (2023a), I2-GNN Huang et al. (2023)
  • K-WL/K-FWL GNNs: PPGN Maron et al. (2019), 2-DRFWL(2) Zhou et al. (2023), 3-DRFWL(2) Zhou et al. (2023)
  • K-FWL+SubgraphGNNs: N2GNN Feng et al. (2023)
  • GD-WL GNNs Graphormer Ying et al. (2021), EPNN Zhang et al. (2024b), CKGConv Ma et al. (2024), GRIT Ma et al. (2023).
For OGB-LSC Hu et al. (2021):
MPNNs (GCN Kipf and Welling (2017), GIN Xu et al. (2019) with/without virtual nodes) as well as several Graph Transformers (GRPE Park et al. (2022), Graphormer Ying et al. (2021), TokenGT Kim et al. (2022) and GraphGPS Rampášek et al. (2022)).
@@ B.1 Sensitivity Study: Impact of the Number of Bases $S$ in SPE on Empirical Expressivity
Table 11: Sensitivity study on $S$ in Sinusoidal encoding for RRWP on CFI graphs from BREC Wang and Zhang (2024).
@@ B.2 Case Study Comparing BN, RMSN and AdaRMSN
Figure 5: (Case Study of AdaRMSN) Visualization of Input and Pred data points [(1) Input; (2) Predictions w/ BN; (3) Predictions w/ RMSN; (4) Predictions w/ AdaRMSN]. RMSN is ineffective in preserving magnitude information, whereas both BN and AdaRMSN successfully maintain the crucial magnitude information of the data point
@@
To better understand the advantages of AdaRMSN over BatchNorm (BN), we conducted a sensitivity study comparing normalization techniques w.r.t. batch sizes, based on PPGT on the ZINC dataset The study (as shown in Fig. 6) showcases the better stability of AdaRMSN compared to the choice of BN used in many previous GTs Kreuzer et al. (2021); Rampášek et al. (2022); Ma et al. (2023).
@@
B.4 Runtime and GPU Consumption
We report basic statistics on the runtime and GPU memory consumption of PPGT, in comparison with GRIT (see Table 12). The results suggest that, even without efficiency-driven techniques, adopting a plain Transformer architecture—rather than more complex graph-specific Transformer designs used in prior work—can yield improved runtime and reduced GPU memory consumption.
@@ C.1 Attention in Graph Transformers
In SAN Kreuzer et al. (2021) and GRIT Ma et al. (2023), the attention mechanism is implemented with the sparse operations in PyTorch Geometric due to the complicated attention mechanism. For PPGTs, we provide two versions, one based on the sparse operations and another one based on the dense operations with padding. The latter one is typically faster for larger-scale graphs but consumes more memory.
@@ C.2 Initializations of Parameters
Following the most recent plain Transformers, we initialize the weights of linear layers in the backbone as well as the prediction heads with truncated normalization with standard deviation $\sigma=0.02$.
For the stems, we utilize Kaiming uniform initialization He et al. (2015) with $a=0$ for hidden layers in MLPs and $a=1$ for output layers in MLPs or standalone linear layers.
@@ C.3 Injection of Degree and Graph Order
For superpixel datasets (CIFAR10, MNIST), we also include the location of pixel into the graph PE.
@@ D.1 GRIT’s Attention
The attention mechanism in GRIT Ma et al. (2023) adopts the conditional MLP Perez et al. (2018), involves linear projections, elementwise multiplications, and an uncommon non-linearity in the form of a signed-square-root:
@@
where $\sigma$ is a non-linear activation (ReLU by default); $\mathbf{W}_{\text{Q}},\mathbf{W}_{\text{K}},\mathbf{W}_{\text{Ew}},\mathbf{W}_{\text{Eb}}\in\mathbb{R}^{d^{\prime}\times d}$, $\mathbf{W}_{\text{A}}\in\mathbb{R}^{1\times d^{\prime}}$, $\mathbf{W}_{\text{V}}\in\mathbb{R}^{d\times d}$ and $\mathbf{W}_{\text{Ev}}\in\mathbb{R}^{d\times d^{\prime}}$ are learnable weight matrices; $\odot$ indicates elementwise multiplication; and $\rho(\mathbf{x}):=(\text{ReLU}(\mathbf{x}))^{1/2}-(\text{ReLU}(-\mathbf{x}))^{1/2}$ is the signed-square-root. Note that $\mathbf{e}_{i,j}$ here corresponds to $\mathbf{p}_{ij}$ in our PPGT, but it requires updating within the attention mechanism. According to Ma et al. (2023), the signed-square-root $\rho$ is necessary to stabilize the training.